OpenStudy (anonymous):

the ratio of orange juice to pineapple juice in tropical treat punch is 4 to 3 Bill has 64 oz of orange juice. how much pineapple juice does he need?

OpenStudy (johnweldon1993):

orange to pinapple is 4 to 3 64 of orange = how much pineapple? 4/3 = 64/x multiply both sides by 3 4 = 192/x x = 192/4 x = 48

OpenStudy (anonymous):

so 4/3 =6/x when x = 3?

OpenStudy (johnweldon1993):

huh? the solution i just wrote out...would be saying that to satisfy the 4:3 amount....for 64oz of oj...there is 48 oz of pineapple juice

OpenStudy (anonymous):

ok now i'm getting it thanks

OpenStudy (johnweldon1993):

oh okay...i thought perhaps you were bringing in another question...lol....but okay...good im glad you understand it :)